• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 00:32:25 on November30,2007

    Entry number 217167

    keyword=polarity change / beam update

    The techs have already switched the polarity of the L-HRS dipole
    to positive for positrons. Now we switch the polarity of Q1,Q2,Q3.
    We also cycled Q2, Q3 with regulator off. Then turned on regulator
    to arrive at the required P0. Fig 1 shows the magnets now. See
    the red "P" for positive on L-HRS.

    Beam update: Estimate return at 2 am. There's a broken power supply
    for BLM.

    FIGURE 1